i need help on elimination method
please help me to solve this equation
5x-2y=3
2x+5y=4

Answer by mananth(16946)

5 x -2 y = 3 .............1

2 x + 5 y = 4 .............2
Eliminate y
multiply (1)by 5
Multiply (2) by 2
25 x -10 y = 15
4 x + 10 y = 8
Add the two equations
29 x = 23
/ 29
x = 4/5
plug value of x in (1)
5 x -2 y = 3
4 -2 y = 3
-2 y = 3 -4
-2 y = -1
y = 1/2
x=4/5, y= 1/2